OpenStudy (anonymous):

Solve 3/7 + 7/9 . Then use estimation to check for reasonableness. Show your work.

OpenStudy (anonymous):

find the common denominator which is 63 the you get 1 and 13/63

OpenStudy (anonymous):

lcm = 63 = 27/63 + 49/63 = 76/63 = 1 13/63 estimation 3/7 is approx. 0.4 7/9 ........ 0.8 total = 1 .2

OpenStudy (anonymous):

so which one is right